# E-values, Multiple Testing and Beyond

### v1

@@ A.5 Proof of Theorem 5
where to get the second equality we have used the fact that when the $i$th hypothesis is rejected (i.e., $p_{i}\leq T_{\mathrm{BH}}$), $R_{\mathrm{BH}}=R_{\mathrm{BH}}(p_{i}\rightarrow 0)$. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\{p_{1},\cdots,p_{i-1},0,p_{i+1},\cdots,p_{n}\}$. We have
@@
For the BC procedure, let ~$\widetilde{\mathcal{F}}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. Then, we have
@@
$\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{1}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\widetilde{\mathcal{F}}_{i}]\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\right],$
where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because both $\max_{i}T_{\mathrm{BH},i}$ and $T_{\mathrm{BC},i}$ are measurable with respect to ~$\widetilde{\mathcal{F}}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$ and thus satisfies Condition (8).
By Proposition 5 and following the same discussion used to prove (A.2), we obtain
$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}.$
Hence we have
@@ A.6 Proof of Theorem 6
We only present the proof for the case of $G=2$. The arguments can be generalized to the general case without essential difficulty. Let us consider the first group. Following the same discussion in the proof of Theorem 4, we have
$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}}\right]$
$=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}\right]$
$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}\right].$
By Proposition 5 and the discussion in the proof of Theorem 4, we have
$\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}=\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}}.$
Thus,
$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq\max_{i\in\mathcal{G}_{1}}w_{i}\mathbb{E}\left[\frac{n_{1}\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}}\right]\leq n_{1}\max_{i\in\mathcal{G}_{1}}w_{i}.$
Using the same argument for the second group, we obtain
Hence, by (17), we deduce that
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n,$
which completes the proof. ∎
A.7 Proof of Theorem 7
Proof.
We only present the proof for the case of $G=2$. The arguments can be generalized to the general case without essential difficulty. Let us consider the first group. Following the same discussion in the proof of Theorem 4, we have
@@ A.7 Proof of Theorem 7
$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\right].$
By Proposition 5 and the discussion in the proof of Theorem 4, we have
@@ A.8 Proof of Theorem 8
By Proposition  6 and Assumption 2, we have
@@
If $\hat{\varphi}_{i}(1-p_{i})>T_{1,i}$, both sides are equal to 00. If $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}$, we claim that $\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}=\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}$. Indeed, if $\hat{\varphi}_{j}(1-p_{j})>T_{1,i}$ but $\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}$, then we have $T_{1,i}<T_{1,j}$. Hence, $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}<T_{1,j}$. By proposition 6, we have $T_{1,i}=T_{1,j}$, which contradicts with $T_{1,i}<T_{1,j}$. The other direction can be proved similarly.
@@ A.9 Proof of Theorem 9
$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\right.$
×$\qquad\times\left.\frac{n_{1}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}}\right]$
@@
where we have used Assumption 1(C) to obtain the first inequality and the second inequality is due to the fact that
@@
By Proposition 6 and the argument in the proof of Theorem 8, we have
@@
Combing the above results leads to
@@
Appendix B Additional proofs
B.1 Proof of Proposition 1
@@
[…]
@@ B.2 Proof of Proposition 2
Let $V_{i}=\mathbbm{1}\{H_{i}\text{ is rejected}\}$. We have
$\mathrm{FDP}(T_{\mathrm{GBH}})=\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{R(T_{\mathrm{GBH}})\vee 1}=\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\frac{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}{R(T_{\mathrm{GBH}})\vee 1}\leq\alpha\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}.$
Therefore, it suffices to prove
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]\leq 1.$
Observing that, for a given $R$, $T_{\mathrm{GBH}}=T_{\mathrm{GBH}}(R)$ is a deterministic function of $R$, we have:
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]=\sum_{i\in H_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{V_{i}\mathbbm{1}\{R=k\}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\right]=\sum_{i\in H_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{V_{i}\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\right],$
where $R(p_{i}\rightarrow 0)$ is the number of rejections obtained by replacing the p-value $p_{i}$ with 0. To clarify the second equality, note that if $V_{i}=0$, the equation is trivially true. When $V_{i}=1$, setting $p_{i}$ to 00 does not change the number of rejections. Using the assumption that $F_{j}(x)=c_{j}h(x)$, we obtain:
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]$ $=\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{1}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k)\}]\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
$\leq\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{F_{i}(T_{\mathrm{GBH}}(k))}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
$=\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{c_{i}}{\sum_{j=1}^{n}c_{j}}\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
$=\frac{\sum_{i\in\mathcal{H}_{0}}c_{i}}{\sum_{j=1}^{n}c_{j}}\leq 1.$
B.3 Proof of Proposition 3
@@
Recall from the proof of Proposition 2 that $T_{\mathrm{GBH}}=T_{\mathrm{GBH}}(R_{\mathrm{GBH}})$ is a deterministic function of $R_{\mathrm{GBH}}$. Additionally, for any $i\in\mathcal{H}_{0}$ and $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}$, replacing the p-value $p_{i}$ with 0 does not change the number of rejections, i.e., $R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=R_{\mathrm{GBH}}$. Therefore, we can infer that
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}e_{i}\right]=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\mathbb{E}\left[\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}},R_{\mathrm{GBH}}=k\}}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}})}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{n\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k),R_{\mathrm{GBH}}=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
$=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{n\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k),R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{nF_{i}(T_{\mathrm{GBH}}(k))\mathbb{E}[\mathbbm{1}\{R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
$=$ $\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{nc_{i}h(T_{\mathrm{GBH}}(k))}{\sum_{j=1}^{n}c_{j}h(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]$
$=$ $n\frac{\sum_{i\in\mathcal{H}_{0}}c_{i}}{\sum_{j=1}^{n}c_{j}}\leq n,$
which completes the proof. ∎
B.4 Proof of Proposition 4
Write $T=T_{\mathrm{GBC}}$ for the ease of notation. First note that
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1\vee\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(p_{j})\leq T\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1\vee\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(p_{j})\leq T\}}\frac{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]$
$\leq$ $\alpha\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right].$
Hence, we only need to show that
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]\leq 1.$
One approach to prove FDR control is through the construction of a super-martingale and the use of the optional stopping time theorem. Here we employ an alternative argument based on the leave-one-out technique. Let $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$ and $\mathbf{p}_{-i}=(p_{1},\dots,p_{i-1},\tilde{p}_{i},p_{i+1},\dots,p_{n})$. Define $T_{i}=T(\mathbf{p}_{-i})$, where we view $T$ as a function of the p-values. Notice that if $\varphi_{i}(p_{i})\leq T$, then we have
$\varphi_{i}(p_{i})\leq T\leq T_{\mathrm{up}}<\varphi_{i}(0.5),$
which implies that $p_{i}<0.5$ since $\varphi_{i}$ is increasing. Hence, if the $i$th hypothesis is rejected, then $p_{i}=\tilde{p}_{i}$. Thus, $\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}=\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}$, which further implies that
$\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]$ $=\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\right],$
where we use the fact that if $p_{i}<0.5$, then $\varphi_{i}(1-p_{i})\geq\varphi_{i}(0.5)>T_{\mathrm{up}}\geq T_{i}$. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. For $i\in\mathcal{H}_{0}$, we have
$\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]=$ $\mathbb{E}\left[\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\Bigg{|}\mathcal{F}_{i}\right]\right]$
$=$ $\mathbb{E}\left[\frac{1}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\mathbb{E}\left[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}\lvert\mathcal{F}_{i}\right]\lvert\right]$
$\leq$ $\mathbb{E}\left[\frac{1}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\mathbb{E}\left[\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}\lvert\mathcal{F}_{i}\right]\lvert\right]$
$=$ $\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\right],$
where we use the assumption that $p_{i}$ satisfies Condition (8) to get the inequality. By Proposition  6, we have
$\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}=\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}=\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}.$
If $\varphi_{i}(1-p_{i})>T_{i}$, both sides are equal to 00. If $\varphi_{i}(1-p_{i})\leq T_{i}$, we claim that $\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}=\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}$. Indeed, if $\varphi_{j}(1-p_{j})>T_{i}$ but $\varphi_{j}(1-p_{j})\leq T_{j}$, then we have $T_{i}<T_{j}$. Hence, $\varphi_{i}(1-p_{i})\leq T_{i}<T_{j}$. By proposition 6, we have $T_{i}=T_{j}$, which contradicts with the assumption $T_{i}<T_{j}$. The other direction can be proved similarly.
Hence,
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]\leq\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}\right]\leq 1,$ (B.1)
which finishes the proof.
B.5 Proof of Proposition 5
Proof.
Proposition 5 is a special case of Proposition 6 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
B.6 Proof of Proposition 6
Proof.
@@ B.6 Proof of Proposition 6
Write $T=T_{\mathrm{GBC}}$ for the ease of notation. First, given a p-value vector $\mathbf{p}=(p_{1},\cdots,p_{n})$, recall that the threshold $T$ is defined as
@@
Without loss of generality, let us assume $T_{i}\geq T_{j}$. By the assumption that $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, we have $\varphi_{i}(1-p_{i})\leq T_{i}$ and $\varphi_{j}(1-p_{j})\leq T_{i}$. Since $\varphi_{i}$ is an increasing function, we have $\varphi_{i}(1-p_{i})\leq T_{\mathrm{up}}<\varphi_{i}(0.5)$, which implies $0.51-p_{i}<$ 0.5. Thus $\varphi_{i}(p_{i})\geq\varphi_{i}(0.5)>T_{\mathrm{up}}\geq T_{i}$. The same discussion for $p_{j}$ leads to $\varphi_{j}(p_{j})>T_{i}$.

### v2

@@ B.1 Proof of Theorem 2
where to get the second equality, we have used the fact that when the $i$th hypothesis is rejected (i.e., $p_{i}\leq T_{\mathrm{BH}}$), $R_{\mathrm{BH}}=R_{\mathrm{BH}}(p_{i}\rightarrow 0)$. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\{p_{1},\cdots,p_{i-1},0,p_{i+1},\cdots,p_{n}\}$. We have
@@
For the BC procedure, let $\widetilde{\mathcal{F}}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. Then, we have
@@
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{1}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\widetilde{\mathcal{F}}_{i}]\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\right],$
where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because both $\max_{i}T_{\mathrm{BH},i}$ and $T_{\mathrm{BC},i}$ are measurable with respect to $\widetilde{\mathcal{F}}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$ and thus satisfies Condition (6).
By Proposition B.1, we have
$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}.$ (B.1)
If $p_{i}<1-T_{\mathrm{BC},i}$, both sides are equal to 0. If $p_{i}\geq 1-T_{\mathrm{BC},i}$, we claim that $\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}=\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}$. Indeed, if $p_{j}\geq 1-T_{\mathrm{BC},i}$ but $p_{j}<1-T_{\mathrm{BC},j}$, we have $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$. This implies that $\min(p_{i},p_{j})\geq 1-\max\{T_{\mathrm{BC},i},T_{\mathrm{BC},j}\}$. By Proposition B.1, we have $T_{\mathrm{BC},i}=T_{\mathrm{BC},j}$, which contradicts with the fact that $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$. The other direction can be proved similarly. Hence, we have
@@ B.2 Proof of Theorem 1
We only present the proof for the case of $G=2$. The arguments can be generalized to the general case without essential difficulty. Let us consider the first group. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{1,i}$. Then, we have
@@
$=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{1,i}\}|\mathcal{F}_{i}]\right]$
$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\right],$
where (i) we have used the fact that $T_{1}=T_{1,i}$ when $p_{i}\leq T_{1}<0.5$ to get the second equation, (ii) the third equation follows because $T_{1,i}$ are measurable with respect to $\mathcal{F}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ satisfies Condition (6) under the null.
By Proposition B.1 and the proof of (B.1), we have
@@ B.3 Proof of Theorem 3
By Proposition  B.2 and Assumption 2, we have
@@
If $\hat{\varphi}_{i}(1-p_{i})>T_{1,i}$, both sides are equal to 0. If $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}$, we claim that $\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}=\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}$. Indeed, if $\hat{\varphi}_{j}(1-p_{j})>T_{1,i}$ but $\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}$, then we have $T_{1,i}<T_{1,j}$. Hence, $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}<T_{1,j}$. By proposition B.2, we have $T_{1,i}=T_{1,j}$, which contradicts with $T_{1,i}<T_{1,j}$. The other direction can be proved similarly.
@@ B.4 Proof of Theorem 4
$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\right$.\
$\qquad\times\left.\frac{n_{1}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}}\right]$
@@
where we have used Assumption 1(C) to obtain the first inequality, and the second inequality is due to the fact that
@@
By Proposition B.2 and the argument in the proof of Theorem 3, we have
@@
Combining the above results leads to
@@
Appendix C Supplementary Proofs
C.1 Proof of Proposition 1
@@
$\text{FDP}=$ $\sum_{i=1}^{n}\frac{\mathbbm{1}\{ie_{(i)}\geq n/\alpha,H_{(i)}\text{ is under the null}\}}{1\vee\hat{k}}$
@@
Proposition B.1 is a special case of Proposition B.2 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
C.3 Proof of Proposition B.2
@@
Write $T=T_{\mathrm{FBC}}$ for the ease of notation. First, given a p-value vector $\mathbf{p}=(p_{1},\cdots,p_{n})$, recall that the threshold $T$ is defined as
@@
Without loss of generality, let us assume $T_{i}\geq T_{j}$. By the assumption that $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, we have $\varphi_{i}(1-p_{i})\leq T_{i}$ and $\varphi_{j}(1-p_{j})\leq T_{i}$. Since $\varphi_{i}$ is an increasing function, we have $\varphi_{i}(1-p_{i})\leq T_{\mathrm{up}}<\varphi_{i}(0.5)$, which implies $1-p_{i}<0.5$. Thus $\varphi_{i}(p_{i})\geq\varphi_{i}(0.5)>T_{\mathrm{up}}\geq T_{i}$. The same discussion for $p_{j}$ leads to $\varphi_{j}(p_{j})>T_{i}$.